OpenStudy (anonymous):

You pour 1000 cm^3 of syrup on the floor at the rate of 50 cm^3/s. The syrup flows in all direction equally - creating an expanding circular area as viewed from above. The high viscosity of the syrup causes the syrup to maintain a constant depth of 0.25 cm, creating a thin cylinder. At what rate is the radius changing with respect to time when the radius is 20 cm?

OpenStudy (anonymous):

related rates using derivative of cylinder volume formula, me thinks.

OpenStudy (anonymous):

\[\large \frac{dr}{dt}=\frac{\cancel{dV}}{dt} \cdot \frac{dr}{\cancel{dV}}\]

OpenStudy (amriju):

dV/dt=50cc..V=volume of cylinder... so...d(3.14r^2*0.25)/dt=50....V=3.14r^2*h.....h is height thats const... from herefind dr/dt....that will depend on r itself...so u have dr/dt at r =20...

OpenStudy (anonymous):

Volume, V, of cylinder = πr^2h; find dV/dr from this. dV/dt is given.

OpenStudy (baldymcgee6):

V = pi*r^2*d 1000cm^3 = 2*pi*r*dr/dt * 0.25 Solve for dr/dt, then input r = 20

OpenStudy (anonymous):

The 1000cm^3 isn't necessary here since the syrup at r=20 is much less than that.

OpenStudy (anonymous):

Besides, when you differentiate, the volume formula, you do not get "1000cm^3 = 2*pi*r*dr/dt * 0.25" If you solve this for dr/dt, it will be incorrect.

OpenStudy (lgbasallote):

i wonder if the asker was just deleting his replies...or these users were doing monologues

OpenStudy (amriju):

I am always here to show off...lol

OpenStudy (anonymous):

I started answering, then others offered insights, so I offered commentary and additional info and corrections on that.

OpenStudy (anonymous):

OP seems to be asleep, though . ..

OpenStudy (lgbasallote):

you see it as offering insights...i see it as racing...same shiz...

OpenStudy (anonymous):

? What do you mean, @lgbasallote ?

OpenStudy (lgbasallote):

everyone wants to be the right one....everyone wants to get the medal....so they race with others....

OpenStudy (anonymous):

Interesting interpretation. Does your cynicism often get you to the correct explanation?

OpenStudy (lgbasallote):

my cynicism gets me enemies

OpenStudy (anonymous):

Cynicism aside, how would you rate the flow of the conversation in its ability to help the asker get the correct answer?

OpenStudy (lgbasallote):

no idea. i'm still in algebra

OpenStudy (anonymous):

Very well. Thanks for the input.

OpenStudy (anonymous):

good post. I lol'ed.

OpenStudy (fellowroot):

Why do we all want to be right? Its because it makes us better than everyone else. hahahahahaha